# Mitigating Biases of Large Language Models in Stance Detection with Counterfactual Augmented Calibration

### v3

@@ Abstract
Ang Li1,3,4∗, Jingqian Zhao1,3thanks:  The first two authors contribute equally to this work., Bin Liang1,3,4, Lin Gui5, Hui Wang2, Xi Zeng6, Xingwei Liang1,3, Kam-Fai Wong4, and Ruifeng Xu1,2,3thanks:  Corresponding Author 1 Harbin Institute of Technology, Shenzhen, China 2 Peng Cheng Laboratory, Shenzhen, China 3 Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies 4 The Chinese University of Hong Kong, Hong Kong, China  5 King’s College London, UK 6 The 30th Research Institute of China Electronics Technology Group Corporation, Chengdu, China {angli,zhaojingqian}@stu.hit.edu.cn,  xuruifeng@hit.edu.cn
@@ 1 Introduction
Figure 1: An example demonstrates two types of biases encountered by large language models in stance detection tasks (shown at the top and bottom) as well as unbiased stance rationale (depicted in the middle).
Our experiment reveals that LLMs are influenced by two types of bias patterns in stance detection: 1) sentiment-stance spurious correlations, where the sentiment expressed in the sentence misleads the stance judgment towards a particular target, and 2) target preference bias, where LLMs exhibit a certain stance bias towards certain individuals or topics. Figure 1 illustrates examples of the two types of biases, as well as unbiased stance rationale. Furthermore, our results reveal a significant negative correlation between stance bias and stance detection performance, emphasizing the necessity of alleviating bias to improve the effectiveness of LLMs in stance detection.
@@
Therefore, to address the above two limitations, we propose to mitigate biases of LLMs in stance detection with a Counterfactual Augmented Calibration Network, coined as FACTUAL. We establish a trainable calibration network to approximate the inverse projection function of the bias label distribution within LLMs. This calibration network receives samples, along with stance judgments and rationales from LLMs, and generates calibrated stance judgments. We construct counterfactual augmented data against the training data to rectify stance biases. The counterfactual samples are constructed from both causal and non-causal features, which can enhance the calibration network to yield unbiased stances and accomplish out-of-domain generalization. Through counterfactual augmented supervised training, the calibration network can capture biases present in specific samples, thereby performing debiasing. The main contributions of our work are summarized as follows:
@@
3) A series of experiments demonstrate that our FACTUAL can effectively reduce the bias of LLMs in stance detection, improving the performance in both in-target and zero-shot stance detection tasks [^3] The code is available at https://github.com/Leon-Francis/FACTUAL.
@@ 3.2.1 Sentiment-Stance Spurious Correlations
We conducted experiments in various settings: Task-Des used task-related descriptions for stance judgment, CoT-Demo used the task description with 4-shot chain-of-thought demonstration, and Debias-Instruct used the task description indicating that sentiment was spurious cues for stance judgment. Refer to Appendix C for the detailed prompts. The results are shown in Table 1. We can observe that in most cases, there is a negative correlation between bias-SSC and stance detection performance. See further analysis in Appendix B. Moreover, prompt engineering methods proved ineffectual in mitigating this inherent bias.
@@ 3.2.2 Target Preference Bias
Figure 4: The overall architecture of our proposed FACTUAL. (a) and (b) in the counterfactual data generation represent two ways to generate counterfactual augmentation. $X$ donates the text, $T$ donates the target, $H$ donates the features of the interaction of text and target, and $Y$ donates the stance label. $C$ represents confounding factors, which arise from the two types of biases previously analyzed and may distort the stance prediction.
@@ 4.1 Calibration Network
Specifically, as illustrated in Figure 4, we first use the CoT-Demo instruction (refer to Appendix C for the detail) to obtain the stance judgment and rationale from the LLMs. Then, we input the sample, along with this stance judgment and rationale, into our calibration network (using RoBERTa-base in our setup) to obtain the debiased stance output. We train the calibration network using the cross-entropy loss function with ground truth label:
@@
Through supervised training, our calibration network can capture biases present in specific samples, thereby performing debiasing.
@@ 4.2 Counterfactual Data Augmentation
One challenge in supervised training is the limited representation of bias within the overall training set, and the learned bias features are difficult to generalize. To fully leverage our analysis of the existing stance biases in LLMs (Section 3), facilitate the calibration network to learn diverse bias patterns, we generate the Counterfactual Augmented Data (CAD) against the training data.
First, we construct counterfactual samples by perturbing non-causal features, approaching from the perspective of bias causation. As shown in counterfactual data generation (a) in Figure 4, to mitigate the impact of sentiment-Stance spurious correlations, we perturb the text $x_{i}$ against the spurious correlation of sentiment. To mitigate the impact of target preference bias, we perturb the target $t_{i}$ against the target preference. Specifically, we construct an instruction that allows the LLMs to rephrase the original sentence and target with different words and sentiments while ensuring that the semantics and the stance remain unchanged. Refer to Figure 10 for the detailed prompts. This obtains the perturbed text $x_{i}^{*}$ and perturbed target $t_{i}^{*}$. Since we only disturbed confounding, the stance label remains unaffected. We construct cross-entropy loss on non-causal counterfactual augmented data as follows:
@@
Second, we construct counterfactual samples from the perspective of actual stance causality. As shown in counterfactual data generation (b) in Figure 4, we make necessary alterations to text $x_{i}$ to reverse the stance to target $t_{i}$, thereby only perturbing the causal features. Refer to Figure 11 for the detailed prompts. This obtains the perturbed text $\widetilde{x_{i}}$ expressing a reversed stance to target $t_{i}$. We construct cross-entropy loss on causal counterfactual augmented data as follows:
@@ 5.3 Evaluation Metric
Across three datasets, we used the same evaluation metric established by their proposers, which was also adopted by most of the subsequent baselines. We adopt the macro-average of the F1-score as the evaluation metric. For Sem16 and P-Stance, we report $F1=(F_{favor}+F_{against})/2$. For VAST, we report $F1=(F_{favor}+F_{against}+F_{none})/3$. For in-target stance detection, we select each target to divide training, validation, and test sets. In zero-shot stance detection, for Sem16 and P-Stance, we use the leave-one-target-out evaluation setup. For the VAST dataset, we use their original zero-shot dataset settings. We use standard train/validation/test splits for in-target and zero-shot stance detection across the three datasets.
@@ 6.3 Mitigating Biases Effect Analysis
Table 6: Results of Bias-SSC and Bias-TPB in in-target stance detection on Sem16, P-Stance, and zero-shot stance detection on VAST. The best scores are in bold.
@@ 6.4 Ablation Study
Table 7: Experimental results of ablation study of in-target stance detection on the Sem16, and zero-shot stance detection on VAST. Results on P-Stance are shown in Table 10. The best scores are in bold.
@@ 7 Conclusion
Our framework involves using GPT-3.5 to generate counterfactual augmented data. As we discussed in Appendix F, these samples may contain errors, but overall are beneficial to the training of our calibration network. The methods of constructing counterfactual augmented data using manual annotation or other methods remain to be explored.
@@ Appendix D Experimental Result of LLMs Bias
Table 10: Experimental results of ablation study of in-target stance detection on the P-Stance dataset. The best scores are in bold.

### v4

@@ Abstract
Ang Li1,3∗, Jingqian Zhao1thanks:  The first two authors contribute equally to this work., Bin Liang1,3,4, Lin Gui5, Hui Wang2, Xi Zeng6, Xingwei Liang1, Kam-Fai Wong3,4, and Ruifeng Xu1,2thanks:  Corresponding Author 1 Harbin Institute of Technology, Shenzhen, China  2 Peng Cheng Laboratory, Shenzhen, China 3 The Chinese University of Hong Kong, Hong Kong, China 4 MoE Key Lab of High Confidence Software Technologies, CUHK, China 5 King’s College London, UK  6 The 30th Research Institute of China Electronics Technology Group Corporation, Chengdu, China {angli,zhaojingqian}@stu.hit.edu.cn,  xuruifeng@hit.edu.cn
@@ 1 Introduction
Figure 1: An example demonstrates two types of biases encountered by large language models in stance detection tasks (shown at the top and bottom) as well as unbiased stance rationale (shown in the middle).
Our experiment identifies two primary biases in LLMs for stance detection: (1) sentiment-stance spurious correlations, where sentiment misleads stance judgment, and (2) target preference bias, where LLMs favor certain individuals or topics. Figure 1 illustrates examples of the two types of biases, as well as unbiased stance rationale. Furthermore, our results reveal a significant negative correlation between stance bias and stance detection performance, emphasizing the necessity of alleviating bias to improve the effectiveness of LLMs in stance detection.
@@
Therefore, to address the above two limitations, we propose to mitigate biases of LLMs in stance detection with a Counterfactual Augmented Calibration Network, coined as FACTUAL. We establish a trainable calibration network to approximate the inverse projection function of the bias label distribution within LLMs. This calibration network takes samples as input, including stance judgments and rationales from LLMs, and generates calibrated stance judgments. We construct counterfactual augmented data against the training data to rectify stance biases. The counterfactual samples are constructed from both causal and non-causal features, which can enhance the calibration network to yield unbiased stances and accomplish out-of-domain generalization. Through counterfactual augmented supervised training, the calibration network can capture biases present in specific samples, thereby performing debiasing. The main contributions of our work are summarized as follows:
@@
3) A series of experiments demonstrate that our FACTUAL can effectively reduce the bias of LLMs in stance detection, improving the performance in both in-target and zero-shot stance detection tasks [^3] The code is available at https://github.com/Leon-Francis/FACTUAL..
@@ 3.2.1 Sentiment-Stance Spurious Correlations
We conducted experiments in various settings: Task-Des used task-related descriptions for stance judgment [^4] The results presented in Figures 2 and 3 are obtained based on Task-Des., CoT-Demo used the task description with 4-shot chain-of-thought demonstration, and Debias-Instruct used the task description indicating that sentiment was spurious cues for stance judgment. Refer to Appendix C for the detailed prompts. The results are shown in Table 1. We can observe that in most cases, there is a negative correlation between bias-SSC and stance detection performance. See further analysis in Appendix B. Moreover, prompt engineering methods proved ineffectual in mitigating this inherent bias.
@@ 3.2.2 Target Preference Bias
Figure 4: The overall architecture of our proposed FACTUAL. (a) and (b) in the counterfactual data generation represent two ways to generate counterfactual augmentation. $X$ donates the text, $T$ donates the target, $H$ donates the features of the interaction of text and target, and $Y$ donates the stance label. $C$ represents confounding factors, which arise from the two types of biases previously analyzed and may distort the stance prediction. * denotes the perturbation of non-causal features, and $\sim$ denotes the perturbation of causal features.
@@ 4.1 Calibration Network
Specifically, as illustrated in Figure 4, we first use the CoT-Demo instruction (refer to Appendix C for the detail) to obtain the stance judgment and rationale from the LLMs (which correspond to the result of an argmax operation over the $P_{obs}$). Then, we input the sample, along with this stance judgment and rationale, into our calibration network (using RoBERTa-base in our setup) to obtain the debiased stance output. We train the calibration network using the cross-entropy loss function with ground truth label:
@@ 4.2 Counterfactual Data Augmentation
One challenge in supervised training is the limited representation of bias within the overall training set, and the learned bias features are difficult to generalize. To fully leverage our analysis of the existing stance biases in LLMs (Section 3) and facilitate the calibration network in learning diverse bias patterns, we generate non-causal and causal Counterfactual Augmented Data (CAD) based on the training data.
The objective of non-causal counterfactual data augmentation is to explicitly perturb bias-inducing features in the data, enabling the model to identify and mitigate biases. As illustrated in Figure 4 (a), we construct counterfactual samples by modifying non-causal features while preserving stance labels. Specifically, we construct an instruction that allows the LLMs to perturb the sentence and target. To address sentiment-stance spurious correlations, we perturb the text $x_{i}$ by altering sentiment-related expressions while maintaining the stance. Similarly, to counteract target preference bias, we modify the target $t_{i}$ while ensuring the stance remains unchanged. Refer to Figure 10 for the detailed prompts. This obtains the perturbed text $x_{i}^{*}$ and perturbed target $t_{i}^{*}$. Since we only disturbed confounding, the stance label remains unaffected. We construct cross-entropy loss on non-causal counterfactual augmented data as follows:
@@
In contrast, causal counterfactual data augmentation directly manipulates causal features by reversing the stance. This augmentation improves the model’s capacity to focus on causal features, thereby reducing its reliance on spurious correlations that may introduce bias. As illustrated in Figure 4 (b), we make necessary alterations to text $x_{i}$ to reverse the stance to target $t_{i}$, thereby only perturbing the causal features. Refer to Figure 11 for the detailed prompts. This obtains the perturbed text $\widetilde{x_{i}}$ expressing a reversed stance to target $t_{i}$. We construct cross-entropy loss on causal counterfactual augmented data as follows:
@@ 5.3 Evaluation Metric
Across three datasets, we used the same evaluation metric established by their proposers, which was also adopted by most of the subsequent baselines. We adopt the macro-average of the F1-score as the evaluation metric. For Sem16 and P-Stance, we report $F1=(F_{favor}+F_{against})/2$. For VAST, we report $F1=(F_{favor}+F_{against}+F_{none})/3$. In the in-target stance detection setting, the model is trained and tested on the same set of targets. We follow the dataset splits provided by the original dataset publisher to ensure a fair comparison with related baselines. The zero-shot stance detection setting presents the model with unseen targets during testing, requiring it to generalize from known targets to infer stances toward new ones. For Sem16 and P-Stance, we take one target as the test set while splitting the remaining targets into training and validation sets in a 7:1 ratio. For VAST, which is inherently a zero-shot stance detection dataset, we use the original splits provided by the dataset publisher. These settings are consistent with the baselines to ensure fairness in comparison.
@@ 6.3 Mitigating Biases Effect Analysis
Table 6: Results of Bias-SSC and Bias-TPB in in-target stance detection on Sem16, P-Stance, and zero-shot stance detection on VAST. The best scores are in bold. Results with $\star$ denote the significance tests of our FACTUAL over the same type baseline models at p-value < 0.05.
@@ 6.4 Ablation Study
Table 7: Experimental results of ablation study of in-target stance detection on the Sem16, and zero-shot stance detection on VAST. Results on P-Stance are shown in Table 10. The best scores are in bold. Results with $\star$ indicate the significance tests of our FACTUAL over the ablation experiments at p-value < 0.05.
@@ 7 Conclusion
Acknowledgements
This work was partially supported by the National Natural Science Foundation of China 62176076, Natural Science Foundation of Guang Dong 2023A1515012922, the Shenzhen Foundational Research Funding JCYJ20220818102415032, the Major Key Project of PCL2021A06, CIPSC-SMP-ZHIPU Large Model Cross-Disciplinary Fund ZPCG20241119405 and Hong Kong RGC GRF No. 14206324.
@@ Acknowledgements
Our framework involves using GPT-3.5 to generate counterfactual augmented data. As we discussed in Appendix F, these samples may contain errors, but overall are beneficial to the training of our calibration network. The methods of constructing counterfactual augmented data using manual annotation or other methods remain to be explored. While our study primarily focuses on stance biases—particularly sentiment-stance spurious correlations and target preference bias—our approach can be extended to mitigate other forms of bias by designing appropriate counterfactual samples.
@@ Appendix D Experimental Result of LLMs Bias
Table 10: Experimental results of ablation study of in-target stance detection on the P-Stance dataset. The best scores are in bold. Results with $\star$ indicate the significance tests of our FACTUAL over the ablation experiments at p-value < 0.05.